f(x) = -2x
finding the ordered of pairs for this problem.. not understand this to well. (1,-2) is one that i came up with if some one may help me come up with a few more that would be great and explain to me a little how they worked to get there answer. Thanks.

Answer by stanbon(75887)

f(x) = -2x
finding the ordered pairs for this problem.. not understand this to well.
(1,-2) is one that i came up with if some one may help me come up with a few more
---------------------------
x........f(x) = -2x
-------------------------
1........f(1) = -2*1 = -2
---
2........f(2) = -2*2 = -4
---
3........f(3) = -2*3 = -6
---
10.......f(10)= -2*10 = -20
---
etc.
Cheers,
Stan H.